# K-COMP: Retrieval-Augmented Medical Domain Question Answering With Knowledge-Injected Compressor

### v1

@@ Abstract
Retrieval-augmented question answering (QA) integrates external information, and thereby increases the QA accuracy of reader models that lack domain knowledge. However, documents retrieved for closed domains require high expertise, so the reader model may have difficulty fully comprehending the text. Moreover, the retrieved documents contain thousands of tokens, some unrelated to the question. As a result, the documents include some inaccurate information, which could lead the reader model to mistrust the passages and could result in hallucinations. To solve these problems, we propose K-comp (Knowledge-injected COMPressor) which provides the knowledge required to answer correctly. The compressor automatically generates the requisite prior knowledge to facilitate the answering process prior to the compression of retrieved passages. Subsequently, the passages are compressed autoregressively, with the generated knowledge being integrated into the compression process. This process ensures alignment between the question intent and the compressed context. By augmenting this prior knowledge and concise context, the reader models are guided toward relevant answers and trust the context.
@@
Jeonghun Cho1  Gary Geunbae Lee1,2 1Graduate School of Artificial Intelligence 2Department of Computer Science and Engineering Pohang University of Science and Technology, Republic of Korea {jeonghuncho, gblee}@postech.ac.kr
@@ 1 Introduction
Retrieval-augmented question answering (QA) is a task where passages related to a question are appended to the prompt such that a reader model can reference them and the infer correct answer (Ahmad et al., 2019; Guo et al., 2021).
@@
However, several limitations impede retrieval-augmented approaches in closed domains with large language models (LLMs) as readers. First, the documents retrieved for closed domains require domain expertise, so the reader may not trust the whole text. When faced with unfamiliar input, the model exhibits an availability bias towards commonly known knowledge, making it more willing to believe in information they can easily recall (Jin et al., 2024). Also, retrieved passages contain thousands of tokens and are sometimes unrelated to the question, which can cause the language model to distrust the passages, perceive them as irrelevant noise, and generate answers that do not consider them. These problems lead to hallucinations (Ji et al., 2023a), which result in the model generating inaccurate answers or inferring plausible but false responses. Lastly, LLMs are sensitive to the order of retrieved documents and the prompting method. Specifically, LLMs can have difficulty finding the necessary information within lengthy input prompts, especially when key information or correct answer clues are located in the middle of the prompt (Liu et al., 2024; Xu et al., 2024b).
To tackle these issues, we propose K-comp (knowledge-injected compressor). We aim to use an autoregressive LLM as a compressor with the domain knowledge needed to answer the question, and increase the alignment of the retrieved passages with the question intent. Additionally, when the compressor is trained domain-related terms and information, it becomes able to recognize the entities that occur in the question and provide descriptions for them. This process is significant for closed domains that require substantial prior knowledge. For retrieval augmentation, we use a large amount of text from domain-specific sources, including Wikipedia. We exploit the advantages of domain relevance by efficiently reusing it when annotating prior knowledge, not just for retrieval. Furthermore, we use a causal masking objective (Aghajanyan et al., 2022) during the training phase to inject domain knowledge into the compressor.
@@
  • Even without domain knowledge in the reader model, K-comp provides the medical jargon description to answer the question, thereby enabling LLMs with diverse backgrounds to handle medical questions more accurately.
@@ 2 Related Work
Prompt Compression.
Several studies have demonstrated that prompt augmentations effectively enhance the performance of LLMs across various tasks (Liu et al., 2023a; Ram et al., 2023; Ryu et al., 2023; Wang et al., 2024c; Long et al., 2023; Yagnik et al., 2024). Yet, the relevance and reliability of the augmented passages are significant challenges in prompt augmentations. In order to address this issue, recent studies have attempted to directly extract content from ambiguous and lengthy passages. Kim et al. (2024) eliminates irrelevant information while maximizing the extraction of accurate information, whereas Yang et al. (2023) leverages the black-box LLMs by applying a reward-based method during compressor training to generate summaries. RECOMP (Xu et al., 2024a) selects and augments the summary with the highest end-task performance by using prompts in which non-essential summaries are set to empty strings if necessary. LLMLingua (Jiang et al., 2023a) dynamically assigns different compression rates to various components within the prompt. In contrast, K-comp focuses on the keywords needed to answer the question, emphasizing the alignment between the compressed context and the question.
@@ 3 Causal Knowledge Injection
Causal models trained using autoregressive language modeling rely exclusively on the context to the left of the generated tokens to predict subsequent tokens (Brown et al., 2020). This attribute confers an advantage in causally generating entire documents, such as text generation. However, these models show limited proficiency in tasks that require an understanding of post-positional relationships for span infilling. Conversely, masked language models excel at predicting masked spans by referencing attention scores from tokens located both anteriorly and posteriorly. Nonetheless, their training objective is limited to decoding only short segments of the passages (Devlin et al., 2019; Joshi et al., 2020).
@@ 4.3 K-comp
This approach facilitates the incorporation of descriptions into the prompt for the reader model and ensures that the generated entities and their descriptions are regressively encoded. As a result, a summary is generated in a causal manner with attention on the entities within the question and their related knowledge, thereby composing a summary centered on these domain entities.
@@
Inference.
At inference time, documents are retrieved in advance to construct the compressor input batch $\{\mathcal{P},q_{m}\}$. This enables the sequential autoregressive generation of the entities and descriptions from the question until the <eod> token is produced. The overall context, including entities and descriptions, is then considered, and a summary that aligns more closely with the question is generated. This process ultimately constructs the input prompt for the reader model, ensuring a reliable response to the question.
@@ 5.1 Settings
Evaluation Metrics.
Since all datasets consist of long-form answers, we use the trained model to evaluate answers. We quantify the relevance of answers by using BertScore (Zhang* et al., 2020), which evaluates the similarity between two sentences by exploiting the contextual embeddings of the encoder. We also use UniEval (Zhong et al., 2022), which is a multi-dimensional evaluation metric that has high correlation and similarity with human judgment. We explicitly assess the factual consistency between generated and gold answers.
@@ 5.2 Results
Overall Performance.
Table 1 shows the main results of K-comp compared to the baselines across various reader LLMs. Overall, compression methods are effective. Chunking snippets for retrieval is inherently imperfect, making the Top-1 and Top-5 passages suboptimal. For MedAlpaca, which has the smallest context window size of 2048 among the reader models, answer accuracy declines significantly with Top-5 passages input due to the limited window size. Consequently, a reprocessing stage, such as compression, is required to improve the quality of chunked text and enable the reader model to reference it appropriately. Among baselines, LLMLingua lags behind other baselines trained in the medical domain due to its query-agnostic compression approach. We also observe different results depending on the model size. Larger models are less dynamic in their response, relying more on their internal knowledge and less on the prompt variations. As can be seen in the case study (§6.3) and Table 14, larger models are capable of providing reasonable responses to questions even when presented with noisy input prompts. In other words, the result demonstrates that not only small models but also large models place greater trust in the description and concise text provided by K-comp than in other baselines.
To emphasize the importance of entity and description, we analyze a scenario where K-comp infers normally, but only appends the summary to the reader prompt (Table 2). $-Prior$ is comparable to the baseline fine-tuned for summarization tasks. Even so, it is clear that providing the reader model with prior knowledge significantly improves the accuracy of the final responses compared to FineTune. An interesting observation is that general-purpose models exhibit a more significant influence of information related to medical jargon compared to medical LLMs. Medical LLMs seem to treat knowledge about entities as noisy input, resulting in conflicts with their internal knowledge. Still, these analyses are confined to the QA accuracy of reader LLMs, as they are affected by changes in the components that make up the prompt. The following sections will discuss the relevance and alignment of the summary.
@@ 6.1 Reranking Preference
Table 4: Case study. We provide the passages used to augment the reader prompt and the answers. Red texts highlight the medical jargon within the question. The complete prompt can be found in Table 14
@@ 6.4 GPT-4o Evaluation
Query-Based.
SPLADE (Lassance and Clinchant, 2022) is a lexical-based retriever that reweights each term by emphasizing important terms associated with the query. The most informative top-1 passages among the top-5 retrieved passages are extracted and assumed to be compressed context. GPT-4o-mini 5 is the result of a summary generated using prompt 9, which is the same prompt used when synthesizing the RECOMP train data. The rationale for K-comp being rated higher than RECOMP is that even when trained on that dataset, it produces a summary that lacks specialization, which is similar to the output of FineTune in case study. By using a seq2seq model rather than token-level pruning such as LLMLingua, it allows the delivery of complete sentences to the reader LLM in QA tasks. This prevents the LLM from perceiving excessive noise, thereby ensuring relatively strong end-task performance among the baselines (Table 1). However, due to the maximum token limitation, the generated text is shorter and contains less information in the context, which results in lower evaluation scores for summary quality.
@@
A noteworthy point is the comparison result with query-based baselines. As can be seen in Table 8 and 9, the instructions used for synthesizing summaries in RECOMP and K-comp were as follows: "Compress … used to answer the question" and "Extract the content about the entity" respectively. By focusing on the entities, the objective of our training approach is to provide a concise context of the medical terminology that has been requested in the question. As a consequence, K-comp achieves a win rate that is similar to, and even exceeds, the results obtained in the training datasets. This presents a limitation of previous methods, which are unable to maintain their performance levels due to their reliance on the distribution of training data. On the other side, our training approach focuses on medical jargon, which is advantageous because the medical terminology remains consistent even when the data changes. Therefore, our causal knowledge injection substantially contributes to improving performance in data-scarce, closed-domain settings.
NLG metric.
@@
Table 5: Results evaluated using the G-Eval-4 metric (Liu et al., 2023b). The prompt used is presented in Table 13
The reliability of the methodology has been established through rigorous validation using the BertScore and UniEval metrics. To further reinforce its consistency, we employ G-Eval (Liu et al., 2023b), a state-of-the-art evaluation metric, to conduct a comprehensive supplementary assessment using a robust commercial LLM. In the case of G-Eval-4, GPT-4 is sampled 20 times, with the resulting average score used to minimize the impact of any potential variability. Given the extensive evaluation of our methodology across diverse datasets and a broad spectrum of models, conducting a full-scale assessment with G-Eval would entail substantial computational costs. Consequently, 1k data points per dataset were randomly selected as the test data. The results are reported in Table 5. Regarding performance, K-comp consistently exhibits superior performance in comparison to other baselines.
An interesting observation is that our results demonstrate a trend similar to that of UniEval. According to the G-Eval study, UniEval exhibited a higher correlation with human performance in summary, dialogue generation, and consistency evaluation than all baselines except G-Eval-4 (including G-Eval-3.5 with GPT-3.5 on occasion). This trend is also reflected in our findings. In Table 1, when the reader model is GPT-4o, it is evident that the answers generated by augmenting the top-5 documents have the highest scores, a phenomenon that is also observed in G-Eval. Furthermore, in the MedQuAD dataset, when the Llama-3-70B model is used as the reader, Uni-Eval exhibits a preference for RECOMP, which aligns with the corresponding G-Eval results. Similarly, in other cases, both metrics indicate a preference for K-comp. These findings suggest that UniEval effectively mirrors G-Eval, despite being a relatively older metric. Therefore, although human evaluation is not conducted in our study, the evidence suggests that our methodology is likely to correlate strongly with human judgments.
@@ 7 Conclusion
In this paper, we have proposed a novel method to improve retrieval augmented QA by compressing retrieved documents focused on the questions. We have devised a comprehensive scheme for identifying medical entities and automatically generating prior knowledge. This is followed by the extension of training and inference methods, which enable the autoregressive generation of summaries that incorporate domain knowledge while considering the context causally. Furthermore, we proved that this approach is practical even when applied to unseen evaluation, which represents a novel contribution in closed domains where data is scarce.
@@
In our research, we employed publicly available datasets, including MedQuAD, MASH-QA, BioASQ and MEDIQA. When synthesizing ground-truth summaries, we ensure that no personally identifiable information is used and that all data are anonymized. Our methodology is still in its early stages and is not yet suitable for direct practical use in medical domains, where reliability and accuracy are of utmost importance. In particular, hallucination can critically affect patient care and clinical decision-making. Therefore, our methodology is considered to mitigate hallucination by emphasizing domain knowledge in biomedical QA research, rather than substituting professional medical judgment, thus posing no risk of harm.
Acknowledgements
@@ A.3 Implementation Details
Appendix B Analysis of Query-Agnostic Section
Figure 3 compares our approach to previous studies that compress prompts without relying on the query, as illustrated in the top-left panel. GPT-4o-mini is the result of summarizing only passages without questions. We referred to the prompts used when training RECOMP, which can be seen in Table 10. Selective-Context (SC) (Li et al., 2023) is a method for the removal of contexts with low self-information at the token level. SC was compressed with Llama-2-7B (Touvron et al., 2023) as that used for LLMLingua. Both methods compress prompts by considering token-level dependencies, so the resulting sentences are incomplete when decoding. As can be seen in Table 18, LLMLingua generates sequences in which words and symbols are merely listed. At the same time, SC produces fragments due to token-level pruning, which also results in incomplete sentences. Consequently, although embedding-based metrics such as BertScore (Zhang* et al., 2020) or lexical-based metrics like ROUGE (Lin, 2004) may exhibit robust performance, they are limited in qualitative evaluations. Moreover, the methodology of compressing passages without referencing the query is unsuitable for retrieval-augmented QA tasks. Thus, both GPT-4o-mini and K-comp outperform the aforementioned baselines. In addition, it is expected that K-comp, by referencing masked questions, provides more relevant and specialized context compared to GPT-4o-mini, which does not reference the questions.

### v2

@@ Abstract
Retrieval-augmented question answering (QA) integrates external information and thereby increases the QA accuracy of reader models that lack domain knowledge. However, documents retrieved for closed domains require high expertise, so the reader model may have difficulty fully comprehending the text. Moreover, the retrieved documents contain thousands of tokens, some unrelated to the question. As a result, the documents include some inaccurate information, which could lead the reader model to mistrust the passages and could result in hallucinations. To solve these problems, we propose K-comp (Knowledge-injected compressor) which provides the knowledge required to answer correctly. The compressor automatically generates the prior knowledge necessary to facilitate the answer process prior to compression of the retrieved passages. Subsequently, the passages are compressed autoregressively, with the generated knowledge being integrated into the compression process. This process ensures alignment between the question intent and the compressed context. By augmenting this prior knowledge and concise context, the reader models are guided toward relevant answers and trust the context.
@@
Jeonghun Cho POSTECH GSAI jeonghuncho@postech.ac.kr                        Gary Geunbae Lee POSTECH GSAI POSTECH CSE gblee@postech.ac.kr
@@ 1 Introduction
Retrieval-augmented question answering (QA) is a task where passages related to a question are appended to the prompt such that a reader model can reference them and infer the correct answer (Ahmad et al., 2019; Guo et al., 2021).
@@
However, several limitations impede retrieval-augmented approaches in closed domains with large language models (LLMs) as readers. First, the documents retrieved for closed domains require domain expertise, so the reader may not trust the whole text. When faced with unfamiliar input, the model exhibits an availability bias toward commonly known knowledge, making it more willing to believe in information it can easily recall (Jin et al., 2024). Also, retrieved passages contain thousands of tokens and are sometimes unrelated to the question. This can cause the language model to distrust the passages, perceive them as irrelevant noise, and generate answers that do not consider them. These problems lead to hallucinations (Ji et al., 2023a), which result in the model generating inaccurate answers or inferring plausible but false responses. Lastly, LLMs are sensitive to the order of retrieved documents and the prompting method. Specifically, LLMs can have difficulty finding the necessary information within lengthy input prompts, especially when key information or correct answer clues are located in the middle of the prompt (Liu et al., 2024; Xu et al., 2024b).
To address these issues, we propose K-comp (knowledge-injected compressor). We aim to use an autoregressive LLM as a compressor with the domain knowledge needed to answer the question, and increase the alignment of the retrieved passages with the question intent. Additionally, when the compressor is trained in domain-related terms and information, it becomes able to recognize the entities that occur in the question and provide descriptions for them. This process is significant for closed domains that require substantial prior knowledge. For retrieval augmentation, we use a large amount of text from domain-specific sources, including Wikipedia. We exploit the advantages of domain relevance by efficiently reusing it when annotating prior knowledge, not just for retrieval. Furthermore, we use a causal masking objective (Aghajanyan et al., 2022) during the training phase to inject domain knowledge into the compressor.
@@
  • Even without domain knowledge in the reader model, K-comp provides the description of the medical jargon to answer the question, thereby enabling LLMs with diverse backgrounds to handle medical questions more accurately.
@@ 2 Related Work
Prompt Compression
Several studies have demonstrated that prompt augmentations effectively enhance the performance of LLMs across various tasks (Liu et al., 2023a; Ram et al., 2023; Ryu et al., 2023; Wang et al., 2024c; Long et al., 2023; Yagnik et al., 2024). Yet, the relevance and reliability of the augmented passages are significant challenges in prompt augmentations. In order to address this issue, recent studies have attempted to extract content from ambiguous and lengthy passages directly. Kim et al. (2024) eliminates irrelevant information while maximizing the extraction of accurate information, whereas Yang et al. (2023) leverages the black-box LLMs by applying a reward-based method during compressor training to generate summaries. RECOMP (Xu et al., 2024a) selects and augments the summary with the highest end-task performance by using prompts in which non-essential summaries are set to empty strings if necessary. LLMLingua (Jiang et al., 2023a) dynamically assigns different compression rates to various components within the prompt. In contrast, K-comp focuses on the keywords needed to answer the question, emphasizing the alignment between the compressed context and the question.
@@ 3 Causal Knowledge Injection
Causal models trained using autoregressive language modeling rely exclusively on the context to the left of generated tokens to predict subsequent tokens (Brown et al., 2020). This attribute confers an advantage in causally generating entire documents, such as text generation. However, these models show limited proficiency in tasks that require an understanding of post-positional relationships for span infilling. Conversely, masked language models excel at predicting masked spans by referencing attention scores from tokens located both anteriorly and posteriorly. Nonetheless, their training objective is limited to decoding only short segments of passages (Devlin et al., 2019; Joshi et al., 2020).
@@ 4.3 K-comp
This approach facilitates the incorporation of descriptions into the prompt for the reader model and ensures that the generated entities and their descriptions are regressively encoded. As a result, a summary is generated in a causal manner with attention to the entities within the question and their related knowledge, thereby composing a summary centered on these domain entities.
@@
Inference
At inference time, documents are retrieved in advance to construct the compressor input batch $\{\mathcal{P},q_{m}\}$. This enables the sequential autoregressive generation of entities and descriptions from the question until the <eod> token is produced. The overall context, including entities and descriptions, is then considered, and a summary that aligns more closely with the question is generated. This process ultimately constructs the input prompt for the reader model, ensuring a reliable response to the question.
@@ 5.1 Settings
Evaluation Metrics
Since all datasets consist of long-form answers, we use the trained model to evaluate the answers. We quantify the relevance of answers using BertScore (Zhang* et al., 2020), which evaluates the similarity between two sentences by exploiting the contextual embeddings of the encoder. We also use UniEval (Zhong et al., 2022), which is a multi-dimensional evaluation metric that has high correlation and similarity with human judgment. We explicitly assess the factual consistency between generated and gold answers.
@@ 5.2 Results
Overall Performance
Table 1 shows the main results of K-comp compared to the baselines across various reader LLMs. Overall, compression methods are effective. Chunking snippets for retrieval is inherently imperfect, making the Top-1 and Top-5 passages suboptimal. For MedAlpaca, which has the smallest context window size of 2048 among the reader models, the answer accuracy declines significantly with Top-5 passages input due to the limited window size. Consequently, a reprocessing stage, such as compression, is required to improve the quality of chunked text and enable the reader model to reference it appropriately. Among the baselines, LLMLingua lags behind other baselines trained in the medical domain due to its query-agnostic compression approach. We also observe different results depending on the model size. Larger models are less dynamic in their response, relying more on their internal knowledge and less on the prompt variations. As can be seen in the case study (§6.3) and Table 14, larger models are capable of providing reasonable responses to questions even when presented with noisy input prompts. In other words, the result demonstrates that not only small models but also large models place greater trust in the description and concise text provided by K-comp than in other baselines.
To emphasize the importance of entity and description, we analyze a scenario where K-comp infers normally but only appends the summary to the reader prompt (Table 2). $-Prior$ is comparable to the baseline fine-tuned for summarization tasks. Even so, it is clear that providing the reader model with prior knowledge significantly improves the accuracy of the final responses compared to FineTune. An interesting observation is that general-purpose models exhibit a more significant influence of information related to medical jargon compared to medical LLMs. Medical LLMs seem to treat knowledge about entities as noisy input, resulting in conflicts with their internal knowledge. Still, these analyses are confined to the QA accuracy of reader LLMs, as they are affected by changes in the components that make up the prompt. The following sections will discuss the relevance and alignment of the summary.
@@ 6.1 Reranking Preference
Table 4: Case study. We provide the passages used to augment the reader’s prompt and the answers. Red texts highlight the medical jargon within the question. The complete prompt can be found in Table 14.
@@ 6.4 GPT-4o Evaluation
Query-Based
SPLADE (Lassance and Clinchant, 2022) is a lexical-based retriever that reweights each term by emphasizing important terms associated with the query. The most informative top-1 passages among the top-5 retrieved passages are extracted and assumed to be compressed context. GPT-4o-mini 5 is the result of a summary generated using prompt 10, which is the same prompt used when synthesizing the RECOMP train data. The rationale for K-comp being rated higher than RECOMP is that even when trained on that dataset, it produces a summary that lacks specialization, which is similar to the output of FineTune in case study. By using a seq2seq model rather than token-level pruning such as LLMLingua, it allows the delivery of complete sentences to the reader LLM in QA tasks. This prevents the LLM from perceiving excessive noise, thereby ensuring relatively strong end-task performance among the baselines (Table 1). However, due to the maximum token limitation, the generated text is shorter and contains less information in the context, which results in lower evaluation scores for summary quality.
@@
A noteworthy point is the comparison result with query-based baselines. As can be seen in Table 8 and 10, the instructions used for synthesizing summaries in RECOMP and K-comp were as follows: "Compress … used to answer the question" and "Extract the content about the entity", respectively. By focusing on the entities, the objective of our training approach is to provide a concise context of the medical terminology that has been requested in the question. As a consequence, K-comp achieves a win rate that is similar to, and even exceeds, the results obtained in the training datasets. This presents a limitation of previous methods, which are unable to maintain their performance levels due to their reliance on the distribution of training data. On the other side, our training approach focuses on medical jargon, which is advantageous because the medical terminology remains consistent even when the data changes. Therefore, our causal knowledge injection substantially contributes to improving performance in data-scarce, closed-domain settings.
Additional NLG Evaluation
@@
Table 5: Results evaluated using the G-Eval-4 metric (Liu et al., 2023b). The prompt used is presented in Table 13.
The validity of our method has been established through rigorous empirical validation employing the BertScore and UniEval evaluation metrics. To further enhance methodological consistency and ensure a comprehensive evaluation, we have incorporated G-Eval (Liu et al., 2023b), a cutting-edge assessment metric, to perform an extensive supplementary analysis. This evaluation is conducted leveraging a highly advanced commercial large language model (LLM), thereby reinforcing the reliability and validity of the proposed approach. For G-Eval-4, GPT-4 is sampled 20 times, with the resulting average score used to minimize the impact of any potential variability. Given the broad scope of our methodological evaluation, which covers diverse datasets and a wide range of models, evaluating a full-scale analysis with G-Eval would be extremely computationally expensive. As a result, 1k data points per dataset were randomly selected as the test data. The results are reported in Table 5. Regarding performance, K-comp consistently exhibits superior performance compared to other baselines.
Interestingly, our results align closely with the trends observed in UniEval. According to the G-Eval study, UniEval exhibited a stronger correlation with human judgments in summary, dialogue generation, and consistency evaluation than all baselines except G-Eval-4. This trend is also reflected in our findings. In Table 1, when GPT-4o is used as the reader model, the responses generated by augmenting the top-5 documents achieve the highest scores, a pattern consistent with G-Eval. Additionally, on the MedQuAD dataset, when Llama-3-70B serves as the reader, UniEval shows a preference for RECOMP, which aligns with the corresponding G-Eval results. Similarly, in other scenarios, both metrics indicate a preference for K-comp. These observations suggest that UniEval effectively mirrors G-Eval despite being a relatively older metric. While our study does not include human evaluation, the strong alignment between UniEval and G-Eval suggests that our methodology is likely to correlate well with human judgments.
@@ 7 Conclusion
In this paper, we have proposed a novel method to improve retrieval-augmented QA by compressing retrieved documents focused on the questions. We have devised a comprehensive scheme for identifying medical entities and automatically generating prior knowledge. This is followed by the extension of training and inference methods, which enable the autoregressive generation of summaries that incorporate domain knowledge while considering the context causally. Furthermore, we proved that this approach is practical even when applied to unseen evaluation, which represents a novel contribution in closed domains where data is scarce.
Acknowledgements
@@
In our research, we employed publicly available datasets, including MedQuAD, MASH-QA, BioASQ, and MEDIQA. When synthesizing ground-truth summaries, we ensure that no personally identifiable information is used and that all data are anonymized. Our methodology is still in its early stages and is not yet suitable for direct practical use in medical domains, where reliability and accuracy are paramount. In particular, hallucination can critically affect patient care and clinical decision-making. Therefore, our methodology is considered to mitigate hallucination by emphasizing domain knowledge in biomedical QA research rather than substituting professional medical judgment, thus posing no risk of harm.
@@ A.3 Implementation Details
Appendix B Analysis of Query-Agnostic
Figure 3 compares our approach to previous studies that compress prompts without relying on the query, as illustrated in the top-left panel. GPT-4o-mini is the result of summarizing only passages without questions. We referred to the prompts used when training RECOMP, which can be seen in Table 10. Selective-Context (SC) (Li et al., 2023) is a method for the removal of contexts with low self-information at the token level. SC was compressed with Llama-2-7B (Touvron et al., 2023) as that used for LLMLingua. Both methods compress prompts by considering token-level dependencies, so the resulting sentences are incomplete when decoding. As can be seen in Table 18, LLMLingua generates sequences in which words and symbols are merely listed. At the same time, SC produces fragments due to token-level pruning, which also results in incomplete sentences. Consequently, although embedding-based metrics such as BertScore (Zhang* et al., 2020) or lexical-based metrics like ROUGE (Lin, 2004) may exhibit robust performance, they are limited in qualitative evaluations. Moreover, the methodology of compressing passages without referencing the query is unsuitable for retrieval-augmented QA tasks. Thus, both GPT-4o-mini and K-comp outperform the aforementioned baselines. In addition, it is expected that K-comp, by referencing masked questions, provides a more relevant and specialized context compared to GPT-4o-mini, which does not reference the questions.